# Generalist World Model Pre-Training for Efficient Reinforcement Learning

### v1

@@ Abstract
Sample-efficient robot learning is a longstanding goal in robotics. Inspired by the success of scaling in vision and language, the robotics community is now investigating large-scale offline datasets for robot learning. However, existing methods often require expert and/or reward-labeled task-specific data, which can be costly and limit their application in practice. In this paper, we consider a more realistic setting where the offline data consists of reward-free and non-expert multi-embodiment offline data. We show that generalist world model pre-training (WPT), together with retrieval-based experience rehearsal and execution guidance, enables efficient reinforcement learning (RL) and fast task adaptation with such non-curated data. In experiments over 72 visuomotor tasks, spanning 6 different embodiments, covering hard exploration, complex dynamics, and various visual properties, WPT achieves 35.65% and 35% higher aggregated score compared to widely used learning-from-scratch baselines, respectively.
Reinforcement Learning, World Models, Model-based Reinforcement Learning, Continuous Control, Machine Learning
@@ 1 Introduction
Figure 1: Aggregate performance on 50 manipulation tasks from Meta-World and 22 locomotion tasks from DMControl with pixel inputs. We pre-train one single world model per benchmark on non-curated offline data, which can be fine-tuned to largely improve RL’s sample efficiency. With 150k online samples, WPT obtains 35.65% and 35% higher normalized scores compared to DrQ v2 and Dreamer v3 using the same samples and matches baselines’ results obtained with higher sample budgets (500k samples for DMControl and 1M samples for Meta-World).
@@
Figure 2: Overview of generalist world model pre-training (WPT). Facing non-curated offline datasets with reward-free, non-expert, and multi-embodiment data, we train a task and embodiment-agnostic world model. The pre-trained world model improves the sample efficiency of RL training over a wide range of tasks.
With the success of scaling laws in vision and language (Brown et al., 2020; He et al., 2022; Kirillov et al., 2023; Touvron et al., 2023), the robotics community has started investigating the use of large-scale offline datasets for robot learning (Brohan et al., 2023b; Walke et al., 2023; Brohan et al., 2023a; O’Neill et al., 2023; Khazatsky et al., 2024; Team et al., 2024). Large-scale offline datasets enable the trained agents to learn a wide range of skills as well as help the agent adapt to new tasks with limited samples (Team et al., 2024).
To leverage offline datasets for decision-making, imitation learning (IL) (Brohan et al., 2023b; Team et al., 2024; Doshi et al., 2024) and offline RL (Levine et al., 2020; Kumar et al., 2020; Fujimoto & Gu, 2021; Kumar et al., 2023) are commonly used. However, these methods have limitations in terms of both data and algorithmic aspects. IL assumes the availability of expert data from human demonstration or specialist RL agents, which usually involves an expensive data collection procedure. Further to this, agents learned by IL are usually sensitive to environmental perturbations such as action delay, lighting conditions, sensor noise, etc (Chae et al., 2022; Zare et al., 2024). Offline RL aims to learn agents from offline data but suffers from training instability, especially for pixel-based observations (Kumar et al., 2020; Lu et al., 2023). Besides the training stability issue, offline RL requires task-specific datasets, labeled with rewards. This poses another challenge for many real-world applications, e.g., robotic manipulation, to utilize offline datasets for new tasks, it is necessary to retrospectively annotate the image-based offline data with rewards, which can be challenging and laborious. Furthermore, fine-tuning an offline RL agent, so-called offline-to-online RL (Nair et al., 2020; Lee et al., 2022; Zhao et al., 2022; Yu & Zhang, 2023; Nakamoto et al., 2024), requires additional tricks to stabilize training and prevent performance collapse caused by the distributional shift in the fine-tuning stage.
In contrast to IL and offline RL, which require curated datasets, many real-world scenarios have data from various sources. To this end, we focus on the more realistic setting where the offline dataset includes non-curated mixed-quality data consisting of reward-free and non-expert data, collected by multiple agents with varying embodiments. This setting has a minimal requirement for offline data, which significantly expands the pool of usable data. The primary research question we aim to answer is:
What is the best way to leverage non-curated offline data for efficient multi-task & multi-embodiment learning?
Definition 1.1 (Non-curated data).
Non-curated data encapsulates all observation-action trajectories, i.e., it can consist of reward-free and non-expert data for any embodiment.
Existing methods for leveraging offline data (e.g. IL and offline RL) typically fail given non-curated data. For example, IL would require manually filtering the dataset to select expert trajectories and offline RL would require retrospective reward labeling. With such non-curated data, a common strategy is to pre-train a visual encoder (Schwarzer et al., 2021; Nair et al., 2022; Parisi et al., 2022; Xiao et al., 2022; Yang & Nachum, 2021; Shang et al., 2024). However, visual pre-training alone fails to fully leverage the rich information that exists in the offline dataset, e.g., dynamics model, informative states, or action prior, etc.
Facing the non-curated offline dataset, training a world model seems a “natural” choice, as it holds the promise of better generalization and better sample efficiency. However, model-based approaches are not commonly used in the offline-to-online setting (Rafailov et al., 2023; Wu et al., 2024). This is partially due to the performance margin between model-based and model-free approaches not being significant (Yu et al., 2020; Kidambi et al., 2020) while model-based approaches involve higher training complexity. In addition, in the typical offline-to-online RL setting, training an offline agent with model-based approaches can be unstable, especially for pixel inputs (Lu et al., 2023), and strong regularization is required during the online training stage (Rafailov et al., 2023).
In this paper, we show that with careful design choices, world model pre-training unlocks more useable offline data and better fits the pre-train and fine-tune paradigm used for scaling. Our key insights are i) instead of pre-training a policy using model-based offline RL as done in previous work (Yu et al., 2020; Kidambi et al., 2020; Rafailov et al., 2023), learning a world model is stable and scalable, ii) unlike previous methods (Hafner et al., 2020, 2021, 2023; Hansen et al., 2024) which learns a world model per task, by simply padding actions to unify the action spaces, a single world model can be pre-trained on the non-curated offline data consisting of data sources from different tasks and different embodiments, iii) via fine-tuning, a pre-trained generalist world model can boost RL’s sample efficiency over a wide range of tasks and embodiments, iv) during fine-tuning, experience rehearsal, and execution guidance can largely improve the performance.
@@
Offline RL Off2On RL RLPD MT IL MT Offline RL WPT (ours)
@@
We extensively evaluate WPT on 72 pixel-based continuous control tasks covering locomotion and manipulation with different action spaces, hard exploration, high dimensions, and complex dynamics. Under a limited sample budget (150k samples), WPT outperforms previous SOTA methods by a decent margin. Specifically, WPT obtains 35.65% and 35% higher normalized scores compared to DrQ v2 and Dreamer v3 under the same sample budget and matches baselines’ results obtained with higher sample budgets (500k samples for DMControl and 1M samples for Meta-World.). For example, WPT enables an agent learning to control an Ant robot to walk forward within 100 trails, while widely used learning-from-scratch baselines need 10-30$\times$ samples. We further demonstrate that, without any modifications, WPT also helps task adaptation, where the agent is required to continually adapt its skill to new tasks. Our results demonstrate strong performance with a large margin over model-free methods when leveraging non-curated offline data. We hope that our results motivate further investigation of model-based approaches for using offline data. In summary, our contributions include:
  • We propose a new and more realistic setting for leveraging offline data where the offline data consists only of reward-free and non-expert multi-embodiment data.
  • We show that generalist world model pre-training along with a series of careful design choices achieves strong performance in 72 visuomotor tasks spanning 6 embodiments.
  • To facilitate further research with non-curated offline data, we open-source our datasets, including 60K trajectories and 10M transitions from the DMControl and Meta-World benchmarks.
2 Related Work
In this section, we review methods which leverage offline data, including pre-training in the context of RL and the so-called generalist agents. See Table 1 for a comparison of what types of data the methods can use.
RL with task-specific offline datasets   Leveraging offline data is a promising direction to improve the sample efficiency in RL. One representative method is Offline RL. Offline RL trains agents from offline data without interacting with the environment. It usually constrains the distance between the learned policy and behavior policies in different ways (Kumar et al., 2019, 2020; Wu et al., 2019; Kostrikov et al., 2021, 2022; Fujimoto & Gu, 2021; Uchendu et al., 2023). However, the performance of the learned policy is highly affected by the quality of offline datasets (Yarats et al., 2022). To continue improving the policy, offline-to-online RL (Nair et al., 2020; Zhao et al., 2022; Lee et al., 2022; Yu & Zhang, 2023; Rafailov et al., 2023) was proposed.
Offline RL and offline-to-online RL commonly suffer from training instability and require additional treatments (Lee et al., 2022; Lu et al., 2023). RLPD (Ball et al., 2023) shows that off-policy RL delivers strong performance by directly concatenating the offline data with online data. However, RLPD still requires reward-labeled task-specific offline data and hasn’t discussed the multi-embodiment cases. Recently, ExPLORe (Li et al., 2023) labels reward-free offline data with approximated upper confidence bound (UCB) of rewards to solve hard exploration tasks. However, ExPLORe obtains good performance when the offline data includes near-expert data of the target tasks, while we consider a more general setting with non-curated offline data.
RL with multi-task offline datasets   Recently, several works have applied offline RL in a multi-task setting (Kalashnikov et al., 2021; Yu et al., 2021; Kumar et al., 2023; Julian et al., 2020; Hansen et al., 2024), but require known rewards.  Georgiev et al. (2024) pre-trains a world model for multi-task RL but is designed for state-based inputs and requires reward-labeled offline data for world model pre-training.
To lift the assumption of known rewards, human labeling (Cabi et al., 2020; Singh et al., 2019), inverse RL (Ng et al., 2000; Abbeel & Ng, 2004) or generative adversarial imitation learning (Ho & Ermon, 2016) can be applied. However, this requires human labor or expert demonstrations. Yu et al. (2022) directly sets zero rewards for unlabelled data, which introduces additional bias.
Another stream of work leverages in-the-wild data for RL training. Most methods focus on representation learning. Stooke et al. (2021); Yang & Nachum (2021); Schwarzer et al. (2021); Shah & Kumar (2021); Yuan et al. (2022); Wang et al. (2022); Parisi et al. (2022); Sun et al. (2023); Ze et al. (2023) investigate different representation learning methods and show promising performance boosts during fine-tuning.
Generalist Agents   RL methods usually perform well on a single task (Vinyals et al., 2019; Andrychowicz et al., 2020), however, this contrasts with humans that can perform multiple tasks well. Recent works have proposed generalist agents that master a diverse set of tasks with a single agent (Reed et al., 2022; Brohan et al., 2023b; Team et al., 2024). These methods typically resort to scalable models and large datasets and are trained via imitation learning (Brohan et al., 2023b, a; O’Neill et al., 2023; Khazatsky et al., 2024). In contrast, we seek to train a generalist world model and use it to boost RL performance for multiple tasks and embodiments.
World models   World models learn to predict future observations or states based on historical information. World models have been widely investigated in online model-based RL (Ha & Schmidhuber, 2018; Hafner et al., 2020; Micheli et al., 2023; Alonso et al., 2024). Recently, the community has started investigating scaling world models (Ha & Schmidhuber, 2018), for example, Hu et al. (2023a); Pearce et al. (2024); Wu et al. (2025); Agarwal et al. (2025) train world model with Diffusion Models or Transformers. However, these models are usually trained on demonstration data. In contrast, we explore the offline-to-online RL setting – closely fitting the successful pre-train and then fine-tune paradigm – and we focus on leveraging reward-free and non-expert data to increase the amount of available data for pre-training.
@@ 3 Methods
In this section, we detail our two-stage approach, which consists of (i) world model pre-training, which learns a generalist (i.e. multi-task & embodiment) world model, given offline data, which rather importantly, includes reward-free and non-expert data, and (ii) RL-based fine-tuning which leverages the pre-trained world model and online interaction in an offline-to-online fashion. See Fig. 2 and Alg. 1.
@@ 3.1 Problem Setup
In this paper, we assume the agent can access a static offline dataset $\mathcal{D}_{\text{off}}$, which consists of trajectories $\{\tau_{\text{off}}^{i}\}_{i=1}^{N_{\text{off}}}$, where $\tau_{\text{off}}^{i}$ includes observations and actions $\{o^{i}_{t},a^{i}_{t}\}_{t=0}^{T}$ collected by unknown behavior policies. This means that (i) rewards $r^{i}_{t}$ are unknown, (ii) $\mathcal{D}_{\text{off}}$ does not necessarily include expert trajectories, and (iii) datasets consist of multi-embodiment data. The agent interacts with the environment to collect labeled trajectories $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$ and stores them in an online dataset $\mathcal{D}_{\text{on}}=\{\tau_{\text{on}}^{i}\}_{i=1}^{N_{\text{on}}}$. The goal of this paper is to learn a high-performance policy from both $\mathcal{D}_{\text{off}}$ and $\mathcal{D}_{\text{on}}$ whilst minimizing the amount of online interaction ($N_{on}$) by leveraging the unlabeled offline data $\mathcal{D}_{\text{off}}$. Removing the requirement of rewards and expert trajectories enables the agent to leverage a large set of diverse datasets.
@@ 3.2 Multi-Embodiment World Model Pre-training
During pre-training, instead of training one model per task – as done in previous work (Hafner et al., 2020, 2021, 2023; Hansen et al., 2024) – we show that a single multi-task & embodiment world model more effectively leverages the offline data. Compared to methods (Yang & Nachum, 2021; Schwarzer et al., 2021; Yuan et al., 2022; Parisi et al., 2022; Ze et al., 2023) that only pre-train representations, pre-training a world model creates a general understanding of the environment, which we show can further boost agents’ performance during fine-tuning.
We adopt a widely used world model in the literature, the recurrent state space model (RSSM) (Hafner et al., 2019) by making several modifications: (i) we remove the task-related losses, (ii) we pad the action with zeros to unify the action dimension of different embodiments, and (iii) we scale the model size to 280M. Given these modifications, we show that RSSMs can successfully learn the dynamics of multiple embodiments and can be fine-tuned for mastering different tasks. Our first stage consists of pre-training the following components:
Sequence model $:h_{t}=f_{\theta}(h_{t-1},z_{t-1},a_{t-1})$
Encoder $:z_{t}\sim q_{\theta}(z_{t}|h_{t},o_{t})$
Dynamics predictor $:\hat{z}_{t}\sim p_{\theta}(z_{t}|h_{t})$
Decoder $:\hat{o}_{t}\sim d_{\theta}(\hat{o}_{t}|h_{t},z_{t}).$
The models $f_{\theta}$, $q_{\theta}$, $p_{\theta}$ and $d_{\theta}$ are optimized jointly by minimizing:
@@
$\mathcal{L}(\theta)=\mathbb{E}_{p_{\theta},q_{\theta}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\theta}(o_{t}|z_{t},h_{t})}_{\text{pixel reconstruction loss}}\\ +\beta\cdot\underbrace{\text{KL}\big{(}q_{\theta}(z_{t}|h_{t},o_{t})\ ||\ p_{\theta}(z_{t}|h_{t})\big{)}}_{\text{latent state consistency loss}}\Big{]}.$ (1)
@@
The first term minimizes the reconstruction error while the second term enables the latent dynamics learning. Note that we removed the task-related objectives, e.g., reward prediction and continue prediction. We note that there is plenty of room to improve the world model’s pre-training, e.g., by leveraging recently developed self-supervised training methods (Eysenbach et al., 2023), or advanced architectures (Vaswani et al., 2017; Gu et al., 2022). These improvements are orthogonal to our method and we leave it as future work.
3.3 RL-based Fine-tuning

### v2

# Efficient Reinforcement Learning by Guiding Generalist World Models with Non-Curated Data

@@ Abstract
Leveraging offline data is a promising way to improve the sample efficiency of online reinforcement learning (RL). This paper expands the pool of usable data for offline-to-online RL by leveraging abundant non-curated data that is reward-free, of mixed quality, and collected across multiple embodiments. Although learning a world model appears promising for utilizing such data, we find that naive fine-tuning fails to accelerate RL training on many tasks. Through careful investigation, we attribute this failure to the distributional shift between offline and online data during fine-tuning. To address this issue and effectively use the offline data, we propose two essential techniques: i) experience rehearsal and ii) execution guidance. With these modifications, the non-curated offline data substantially improves RL’s sample efficiency. Under limited sample budgets, our method achieves a 102.8% relative improvement in aggregate score over learning-from-scratch baselines across 72 visuomotor tasks spanning 6 embodiments. On challenging tasks such as locomotion and robotic manipulation, it outperforms prior methods that utilize offline data by a decent margin.
@@ 1 Introduction
Leveraging offline data offers a promising way to improve the sample efficiency of Reinforcement Learning (RL). Prior work has focused primarily on utilizing curated offline data labeled with rewards [1, 2, 3, 4], which is expensive and laborious to obtain. For instance, leveraging offline datasets for new robotic manipulation tasks requires retrospectively annotating image-based data with rewards. We instead propose expanding the pool of usable offline data by utilizing abundant non-curated data that is reward-free, of mixed quality, and collected across multiple embodiments. This leads to our primary research question:
How can we effectively leverage non-curated offline data for efficient RL?
Typical offline-to-online RL [5, 6, 7, 8, 9] methods fail to utilize non-curated offline data due to their assumption of structured data with rewards. While pre-training visual encoders [10, 11, 12, 13, 14, 15] is a common approach to utilize non-curated offline datasets, it fails to fully leverage the rich information, such as dynamics models, informative states, and action priors. On the other hand, learning world models from offline data appears promising for utilizing the non-curated dataset. However, prior work has explored world model training primarily in settings with known rewards [16, 17, 18] or expert demonstrations [19, 20, 21] or focused solely on visual prediction [22, 23]. Recent approaches [24, 25, 26] have developed novel architectures for world model pre-training using in-the-wild action-free data, but paid limited attention to the fine-tuning process. As a result, despite being trained on massive datasets, these methods show only marginal improvements over training-from-scratch baselines. Additionally, due to the computational costs of RL experiments, previous work evaluated only on a small set of tasks, leaving the effectiveness of the learned world model unclear on broader tasks. In contrast, we extensively evaluate our method on 72 visuomotor control tasks – 12 times more than previous baselines – spanning both locomotion and robotic manipulation, demonstrating consistent improvements over existing approaches.
Through experiments, we observe that naively fine-tuning a world model fails to improve RL’s sample efficiency on many tasks. Through careful investigation, we identify the root cause as a distributional shift between offline and online data during fine-tuning. Specifically, when the offline data distribution does not sufficiently cover the data distribution of downstream tasks, the pre-trained world models struggle to benefit policy learning due to this distribution mismatch. Especially for tasks requiring hard exploration, even when the pre-trained world model performs well on "promising" state-action distributions, the agent often fails to learn because it cannot effectively reach these promising regions. Building on these insights, we propose using offline data in both pre-training and fine-tuning stages, in contrast to previous methods that only consider the offline data for world model pre-training [26, 27, 28]. By this end, we propose Generalist-to-Specialist Adaptation (GSA), which enables generalist world models pre-trained on non-curated data to master a wide range of downstream tasks. GSA has two essential techniques: experience rehearsal and execution guidance. Experience rehearsal mitigates distributional shift by retrieving task-relevant trajectories from offline datasets, while execution guidance promotes exploration by steering the agent toward regions where the world model has high confidence.
@@
Figure 1: Overview of GSA (Generalist-to-Specialist Adaptation). Facing non-curated offline data with reward-free, mixed-quality, and multi-embodiment data, we train a task and embodiment-agnostic world model. Combined with experience rehearsal and execution guidance, the pre-trained world model improves the sample efficiency of RL training over a wide range of tasks.
Equipped with our proposed techniques, GSA demonstrates strong performance across a diverse set of tasks. Specifically, under a limited sample budget (150k samples), GSA achieves a 102.8% relative improvement in aggregate score over learning-from-scratch baselines (DrQ v2 and Dreamer v3), while matching their performance achieved with larger sample budgets. On representative challenging tasks, GSA outperforms baselines that leverage offline data as well as state-of-the-art methods using pre-trained world models by a significant margin. Additionally, without any modifications, we show that GSA improves task adaptation, enabling agents to efficiently adapt their skills to new tasks.
To summarize, our contributions are:
  • (C1)
    We propose a more realistic setting for leveraging offline data that consists of reward-free and mixed-quality multi-embodiment data.
  • (C2)
    We demonstrate that naive world model fine-tuning fails on many tasks due to distributional shift between pre-training and fine-tuning data.
  • (C3)
    We propose two essential techniques, experience rehearsal and execution guidance, to mitigate the distributional gap and encourage exploration.
  • (C4)
    We present GSA, which effectively leverages non-curated offline data and significantly outperforms existing approaches across a diverse set of tasks.
2 Related Work
In this section, we review methods that leverage offline data, including pre-training in the context of RL and world models. See App. C for more related work and Table 1 for an overview comparison.
RL with task-specific offline datasets   Leveraging offline data is a promising direction to improve sample efficiency in RL. One representative approach is offline RL, which trains agents using offline data without environment interaction. These methods typically constrain the distance between learned and behavior policies in different ways [2, 3, 29, 30, 31, 32, 33]. However, policy performance is highly dependent on dataset quality [34]. To enable continued improvement, offline-to-online RL methods [5, 6, 7, 9, 17] were developed, which fine-tune policies trained with offline RL by interacting with environments. MOTO [17] proposes a model-based offline-to-online RL method with reward-labeled data, and requires model-based value expansion, policy regularization, and controlling epistemic uncertainty, while our method leverages reward-free and multi-embodiment data and requires none of the techniques proposed by MOTO.
Typical offline-to-online RL face training instability challenges [5, 16]. To mitigate this issue, RLPD [35] is proposed and demonstrates strong performance by simply concatenating offline and online data, but requires reward-labeled task-specific offline data and does not address multi-embodiment scenarios. ExPLORe [36] labels reward-free offline data using approximated upper confidence bounds (UCB) to solve hard exploration tasks, but relies on near-expert data for the target tasks, while we consider a more general setting with non-curated data.
RL with multi-task offline datasets   Recent work has explored multi-task offline RL [4, 18, 37, 38, 39], but requires known rewards. PWM [40] and TDMPC v2 [18] train world models for multi-task RL but are limited to state-based inputs and reward-labeled data. To handle unknown rewards, approaches like human labeling [41, 42], inverse RL [43, 44], or generative adversarial imitation learning [45] can be used, though these require human labor or expert demonstrations. Yu et al. [46] assigns zero rewards to unlabeled data, which introduces additional bias. Apart from these, there is a line of work that focuses on representation learning from in-the-wild data [10, 12, 14, 27, 47, 48, 49, 50, 51, 52] but fails to utilize rich information in the dataset, such as dynamics.
Recent studies [24, 26, 53] explore world model pre-training with action-free data, focusing on world model architecture design to utilize the action-free data. However, we demonstrate that naive fine-tuning of pre-trained world models fails on challenging tasks, while our method, incorporating experience rehearsal and execution guidance, significantly improves RL performance across 72 tasks.
@@ 2 Related Work
Offline RL Off2On RL RLPD MT Offline RL GSA (ours)
@@ 3 Methods
In this section, we detail our two-stage approach, which consists of (i) world model pre-training, which learns a multi-task & embodiment world model, given offline data, which rather importantly, includes reward-free, mixed-quality data, and (ii) RL-based fine-tuning which leverages the pre-trained world model and online interaction in an offline-to-online fashion. See Fig. 1 for the overview and Alg. 1 for the full algorithm.
@@ 3.1 Problem Setup
In this paper, we assume the agent has access to a non-curated but in-domain offline dataset $\mathcal{D}_{\text{off}}$ with three key characteristics: (i) trajectories lack reward labels $r^{i}_{t}$, (ii) data quality is mixed, and (iii) data comes from multiple embodiments. During fine-tuning, the agent interacts with the environment to collect labeled trajectories $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=1}^{T}$ and stores them in an online dataset $\mathcal{D}_{\text{on}}=\{\tau_{\text{on}}^{i}\}_{i=1}^{N_{\text{on}}}$. Our goal is to learn a high-performance policy by leveraging both $\mathcal{D}_{\text{off}}$ and $\mathcal{D}_{\text{on}}$ while minimizing the required online interactions $N_{\text{on}}$.
@@ 3.2 Multi-Embodiment World Model Pre-training
During pre-training, rather than training separate models per task as in previous work [54, 55, 56], we train one world model per benchmark and demonstrate that a single multi-task & embodiment world model can effectively leverage non-curated data.
Since our primary goal is enabling RL agents to use non-curated offline data rather than proposing a new architecture, we adopt the widely-used recurrent state space model (RSSM) [57] with several modifications: (i) removal of task-related losses, (ii) zero-padding of actions to unify dimensions across embodiments, and (iii) scaling the model to 280M parameters. With these changes, we show that RSSMs can successfully learn the dynamics of multiple embodiments and can be fine-tuned for various tasks.
Our first stage pre-trains the following components:
Sequence model $:h_{t}=f_{\theta}(h_{t-1},z_{t-1},a_{t-1})~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{Encoder}:z_{t}\sim q_{\theta}(z_{t}\,|\,h_{t},o_{t})$
Dynamics predictor $:\hat{z}_{t}\sim p_{\theta}(z_{t}\,|\,h_{t})~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{Decoder}:\hat{o}_{t}\sim d_{\theta}(\hat{o}_{t}\,|\,h_{t},z_{t}).$
The models $f_{\theta}$, $q_{\theta}$, $p_{\theta}$ and $d_{\theta}$ are jointly optimized by minimizing:
@@
$\mathcal{L}(\theta)=\mathbb{E}_{p_{\theta},q_{\theta}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\theta}(o_{t}\,|\,z_{t},h_{t})}_{\text{pixel reconstruction loss}}+\beta\cdot\underbrace{\text{KL}\big{(}q_{\theta}(z_{t}\,|\,h_{t},o_{t})\,\|\,p_{\theta}(z_{t}\,|\,h_{t})\big{)}}_{\text{latent state consistency loss}}\Big{]}.$ (1)
@@
The first term minimizes reconstruction error while the second enables latent dynamics learning. While there is room to improve world model pre-training through recent self-supervised methods [58] or advanced architectures [59, 60], such improvements are orthogonal to our method and left for future work.
3.3 RL-based Fine-Tuning with Rehearsal and Guidance